Brief Exercise 10-07 a1-a2 Blue Spruce Taxi Service uses the units-of-activity method in computing depreciation on its taxicabs. Each cab is expected to be driven 140,000 miles. Taxi no. 10 cost $45,000 and is expected to have a salvage value of $200. Taxi no. 10 is driven 28,000 miles in year 1 and 18,400 miles in year 2. Calculate depreciation cost per mile using unit-of-activity method. (Round answer to 2 decimal places, e.g. 0.50.) Depreciation cost s per mile LINK TO TEXT Compute the depreciation for each year. (Round answers to 2 decimal places, e.g. 2,125.50.) Year 1 Year 2 Depreciation
Answer #1

Depreciation cost per mile=(Cost-Salvage value)/Useful Life

=(45000-200)/140,000=$0.32/mile

Depreciation for:

Year 1=(0.32*28000)=$8960

Year 2=(0.32*18400)=$5888